3. An open rectangular box w/ square ends to hold 6400 cu ft, is to be built at a cost of



Php 75.00 per sq ft. for the base and



Php 25.00 /sq ft for the sides. Find the most economical dimensions.

1
Expert's answer
2021-12-21T02:01:01-0500

Solution:

Let the dimensions be "x,x,y", since it has square ends.

Volume of rectangular box "=6400\ ft^3"

"\Rightarrow x.x.y=6400
\\ \Rightarrow x^2y=6400
\\ \Rightarrow y=\dfrac{6400}{x^2}\ ...(i)"

Total surface area "=S=x.x+2(xy+yx)=x^2+4xy"

So, total cost "=C=(75\times x^2)+(4xy)\times25"

"\Rightarrow C=75 x^2+(100x)(\dfrac{6400}{x^2}) \ [Using\ (i)]
\\=75x^2+640000(\dfrac1x)"

On differentiating w.r.t "x" ,

"C'=150x-\dfrac{640000}{x^2}"

Now put C'=0

"\Rightarrow 150x-\dfrac{640000}{x^2}=0
\\ \Rightarrow 150x^3=640000
\\ \Rightarrow x^3=\dfrac{128000}{3}
\\ \Rightarrow x\approx 34.94"

Again differentiating C' w.r.t x,

"C''=150+\dfrac{128000}{x^3}>0, for\ x=34.94"

Thus, minima exists.

Put the value of x in (i).

"y=\dfrac{6400}{34.94^2}\approx5.24"

Thus, most economical dimensions are 34.94 ft, 34.94 ft and 5.24 ft.